# A Framework for Rapidly Developing and Deploying Protection Against Large Language Model Attacks

### v1

@@ Abstract
I Introduction
@@
Building resilient and adaptive security detection platforms for AI is distinct from traditional detection platforms primarily due to the adaptive nature of attacks and the nondeterministic nature of generative AI applications. Security detection systems are locked in a continuous cybersecurity OODA loop (observe, orient, decide, act) [1] with their attackers. The constant engagement means the focus of a detection platform should shift from expending significant one-off effort to achieve the ”best” detection model at a particular point in time, and instead focus on rapid adaptation and response. As soon as a new detection engine is released, its owner must quickly address any new bypasses that are discovered.
@@
Defensive strategies for LLM deployments have consisted of a mix of technical controls and operational best practices [2] that span a range of effectiveness and cost-efficiency. We use “guardrails” to refer to the deployed detection and blocking layer that mediates model inputs/outputs and enforces safety and security policies. Examples include input and output sanitization (e.g., guardrails that filter out malicious prompts and dangerous commands) [3, 4], continuous monitoring and logging to detect anomalies and signs of attack [5], red teaming and stress testing to discover vulnerabilities and observe model behavior [6], and adversarial training to build out inherent defenses against unauthorized manipulation [7].
@@
LLMs can process vast, unstructured datasets from diverse sources to identify anomalies and patterns. Machine learning-based detection models tend to generalize better than signatures for detections and form the basis of most modern security detection products. They are especially important for protecting LLMs from prompt injection attacks [8]—when an adversary prompts a model to override the developers’ guardrails and executes a malicious instruction. Here, semantic understanding of language is critical and pattern-matching approaches alone are not able to fully capture all the nuanced syntactical variations [9].
@@
However, models often generalize poorly to out-of-distribution (OOD) inputs, including novel attack patterns or techniques against LLMs [10] [11]. As a result, attacks not present in the training distribution (e.g., zero-days) are frequently misclassified. Even when ML models do generalize, their detection accuracy is hampered by operational requirements to deploy these models at very low false positive rates, and further complicated by the low base rate of attacks within legitimate network traffic and user activity [12]. To our knowledge, all current detection systems can be bypassed under some conditions, and there is no known method to prevent adversarial attacks on ML models with 100% certainty [11, 13].
@@
Furthermore, it is hard to guarantee ML model consistency across multiple retraining cycles [14]. A new model deployment could introduce unexpected changes in behavior, such as blocking a previously unflagged input and disrupting customer workflows. Therefore, safely releasing an ML model requires careful observation and gradual customer rollout to gather a statistically significant amount of observations. This validation process can conflict with standard engineering release cycles, which require a much faster quality assurance process.
@@
II Threat Intelligence Operations
@@ II-A Architecture Overview
Central to our Threat Intelligence Platform is a comprehensive and dynamically updated taxonomy [15] that unifies existing AI security frameworks while addressing the unique operational needs of production LLM security. We developed this taxonomy through a systematic review of established security standards, including the OWASP LLM Top 10 [2], MITRE ATLAS [16], and National Institute of Standards and Technology (NIST)’s Adversarial Machine Learning Taxonomy [17]. Unlike existing taxonomies that focus primarily on security or safety concerns, our framework bridges both domains, recognizing that production systems must defend against both malicious attacks and harmful outputs. Each threat category includes detailed subcategories with explicit mappings to both OWASP and MITRE classifications, ensuring compatibility with existing security workflows while providing the granularity needed for detection purposes. This granularity is maintained through an agile update mechanism that incorporates insights from emerging threats and incident analysis, allowing us to quickly integrate new edge cases and novel attack techniques. This taxonomy serves as the basis for our Priority Intelligence Requirements (PIRs), which form a prioritized matrix of variables that guide our intelligence collection efforts [18].
@@
II-B Automated Collection Pipeline
@@
The system employs deduplication mechanisms to prevent redundant processing, and we extract and summarize the text with Jina Reader [19], ensuring consistent and complete content retrieval.
II-C Filtering and Scoring
@@
II-D Threat Analysis and Report Generation
@@
II-E Rapid Signature Development
@@
See more information about safe signature deployment strategy in Section IV-A.
II-F Attack Data Generation
@@
II-G Human Labeling and Feedback
@@ III Data Platform
The Data Platform leverages the distributed data warehouse Snowflake, which prioritizes adaptability over traditional processing paradigms, as shown in Fig. 4. A key architectural decision was choosing between traditional ETL pipelines and our novel warehouse-centric approach. Rather than implementing conventional ETL pipelines with rigid schemas and predetermined processing sequences, which can create brittleness when adapting to new threat types, our design emphasizes flexible data transformation capabilities that can rapidly adapt to evolving requirements. This approach supports dynamic threat response through accelerated operational cycles while maintaining system stability.
@@
(a) Data pipeline architecture for LLM prompt processing, showing iterative enrichment workflow from raw data ingestion through correlation and labeling resulting in a unified table, enabling flexible business logic implementation through SQL views.
(b) Snowflake-based Data Platform architecture illustrating SQL-centric processing augmented with Python UDFs and external service calls, providing extensible primitives for flexible data pipeline operations and cross-source correlation.
@@ III-B Labeling
IV Release Platform
One of the largest challenges in deploying a detection and blocking system like our guardrails is updating its detection components post-release. The primary risk lies in unforeseen shifts in the detection distribution, which could potentially disrupt established customer workflows. We have seen countless examples of security vendors accidentally generating catastrophic levels of false positives that ultimately break critical customer infrastructure [20, 21, 22]. Our Data Platform is designed to minimize the chance of such an event, relying on three major components that need updating. First are signatures, which we can release rapidly to remediate any ongoing issues or concerns. Second are ML detection models; these take longer to update and require additional validation due to their potential to cause unexpected shifts in the detection distribution. Third, the orchestration logic that integrates the previous components into the final detection (Fig. 5).
@@ IV-A Signature Updates
IV-B Model and Logic Updates
@@
V Previous Work
As discussed earlier in this paper, significant research has focused on developing detection methods for LLM attacks and protections against them, consisting of technical controls and operational best practices [2, 4, 5, 6, 7]. But relatively few works address the operational challenges of deploying these defenses in production environments requiring rapid threat response and continuous adaptation. We have not found similar examples of a dedicated threat intelligence function designed specifically for AI-related threats that considers the expanded attack surface and the deepening integrations of generative AI and agentic AI tools in software systems and technology.
Early work in operationalizing LLM security began with the development of Vigil [23] in 2022, which introduced the application of YARA rules to the domain of LLM threat detection. The open-source project demonstrated that pattern-matching techniques could effectively identify prompt injection attempts and other LLM-specific threats while leveraging existing security tooling rather than requiring specialized development of novel tools. Our comprehensive defense platform described in this paper significantly extends this foundational approach by integrating rule-based detection into an enterprise-grade threat intelligence pipeline with automated deployment capabilities, enabling organization-scale protection with the ability to rapidly respond to emerging threats.
Recent advances in prompt injection detection have produced promising approaches, including research that compares the performance of early detection systems for LLM security [24]. Google DeepMind’s CaMeL [25] takes an architectural approach, applying traditional software security principles like control flow integrity to LLM systems. PromptShield [26] achieves detection with low false positive rates through fine-tuning detectors on both conversational and application-structured data, while Attention Tracker [27] introduces training-free detection by analyzing attention patterns within LLMs. However, Hackett et al. [28] demonstrated that many prominent guardrail systems can be bypassed with up to 100% success rates using character injection techniques, highlighting the limitations of point-in-time detection approaches. While comprehensive prevention remains an open problem, our platform reduces the window of vulnerability by translating new threats into protections quickly through rapid signature updates and safe staged deployments.
VI Conclusion

### v2

@@ Abstract
keywords:
Large Language Models, LLM Security, Threat Intelligence, Machine Learning Operations, Rapid Response, Guardrails, Adversarial Attacks, Detection Systems
1 Introduction
@@
Building resilient and adaptive security detection platforms for AI is distinct from traditional detection platforms primarily due to the adaptive nature of attacks and the nondeterministic nature of generative AI applications. Security detection systems are locked in a continuous cybersecurity OODA loop (observe, orient, decide, act) Sager (2015) with their attackers. The constant engagement means the focus of a detection platform should shift from expending significant one-off effort to achieve the ”best” detection model at a particular point in time, and instead focus on rapid adaptation and response. As soon as a new detection engine is released, its owner must quickly address any new bypasses that are discovered.
@@
Defensive strategies for LLM deployments have consisted of a mix of technical controls and operational best practices OWASP Foundation (2025) that span a range of effectiveness and cost-efficiency. We use “guardrails” to refer to the deployed detection and blocking layer that mediates model inputs/outputs and enforces safety and security policies. Examples include input and output sanitization (e.g., guardrails that filter out malicious prompts and dangerous commands) Kumar et al. (2024); Jiang et al. (2023), continuous monitoring and logging to detect anomalies and signs of attack Christiano (2022), red teaming and stress testing to discover vulnerabilities and observe model behavior Bullwinkel et al. (2025), and adversarial training to build out inherent defenses against unauthorized manipulation Jain et al. (2023).
@@
LLMs can process vast, unstructured datasets from diverse sources to identify anomalies and patterns. Machine learning-based detection models tend to generalize better than signatures for detections and form the basis of most modern security detection products. They are especially important for protecting LLMs from prompt injection attacks Liu et al. (2024)—when an adversary prompts a model to override the developers’ guardrails and executes a malicious instruction. Here, semantic understanding of language is critical and pattern-matching approaches alone are not able to fully capture all the nuanced syntactical variations Greshake et al. (2023).
@@
However, models often generalize poorly to out-of-distribution (OOD) inputs, including novel attack patterns or techniques against LLMs Benoit et al. (2024) Peng et al. (2024). As a result, attacks not present in the training distribution (e.g., zero-days) are frequently misclassified. Even when ML models do generalize, their detection accuracy is hampered by operational requirements to deploy these models at very low false positive rates, and further complicated by the low base rate of attacks within legitimate network traffic and user activity Pietraszek and Tanner (2005). To our knowledge, all current detection systems can be bypassed under some conditions, and there is no known method to prevent adversarial attacks on ML models with 100% certainty Peng et al. (2024); Insua et al. (2023).
@@
Furthermore, it is hard to guarantee ML model consistency across multiple retraining cycles Wang et al. (2020). A new model deployment could introduce unexpected changes in behavior, such as blocking a previously unflagged input and disrupting customer workflows. Therefore, safely releasing an ML model requires careful observation and gradual customer rollout to gather a statistically significant amount of observations. This validation process can conflict with standard engineering release cycles, which require a much faster quality assurance process.
@@
2 Threat Intelligence Operations
@@ 2.1 Architecture Overview
Central to our Threat Intelligence Platform is a comprehensive and dynamically updated taxonomy Cisco (2025) that unifies existing AI security frameworks while addressing the unique operational needs of production LLM security. We developed this taxonomy through a systematic review of established security standards, including the OWASP LLM Top 10 OWASP Foundation (2025), MITRE ATLAS MITRE (2024), and National Institute of Standards and Technology (NIST)’s Adversarial Machine Learning Taxonomy Vassilev et al. (2024). Unlike existing taxonomies that focus primarily on security or safety concerns, our framework bridges both domains, recognizing that production systems must defend against both malicious attacks and harmful outputs. Each threat category includes detailed subcategories with explicit mappings to both OWASP and MITRE classifications, ensuring compatibility with existing security workflows while providing the granularity needed for detection purposes. This granularity is maintained through an agile update mechanism that incorporates insights from emerging threats and incident analysis, allowing us to quickly integrate new edge cases and novel attack techniques. This taxonomy serves as the basis for our Priority Intelligence Requirements (PIRs), which form a prioritized matrix of variables that guide our intelligence collection efforts Forum of Incident Response and Security Teams (2025) (FIRST).
@@
2.2 Automated Collection Pipeline
@@
The system employs deduplication mechanisms to prevent redundant processing, and we extract and summarize the text with Jina Reader Jina AI Ltd. (2024), ensuring consistent and complete content retrieval.
2.3 Filtering and Scoring
@@
2.4 Threat Analysis and Report Generation
@@
2.5 Rapid Signature Development
@@
See more information about safe signature deployment strategy in Section 4.1.
2.6 Attack Data Generation
@@
2.7 Human Labeling and Feedback
@@ 3 Data Platform
The Data Platform leverages the distributed data warehouse Snowflake, which prioritizes adaptability over traditional processing paradigms, as shown in Fig. LABEL:fig:platform. A key architectural decision was choosing between traditional ETL pipelines and our novel warehouse-centric approach. Rather than implementing conventional ETL pipelines with rigid schemas and predetermined processing sequences, which can create brittleness when adapting to new threat types, our design emphasizes flexible data transformation capabilities that can rapidly adapt to evolving requirements. This approach supports dynamic threat response through accelerated operational cycles while maintaining system stability.
@@
[…]
@@
[…]
[Snowflake-based Data Platform architecture illustrating SQL-centric processing augmented with Python UDFs and external service calls, providing extensible primitives for flexible data pipeline operations and cross-source correlation.]Refer to caption
@@ 3.2 Labeling
4 Release Platform
One of the largest challenges in deploying a detection and blocking system like our guardrails is updating its detection components post-release. The primary risk lies in unforeseen shifts in the detection distribution, which could potentially disrupt established customer workflows. We have seen countless examples of security vendors accidentally generating catastrophic levels of false positives that ultimately break critical customer infrastructure CrowdStrike (2024); Whittaker (2012); Krebs (2010). Our Data Platform is designed to minimize the chance of such an event, relying on three major components that need updating. First are signatures, which we can release rapidly to remediate any ongoing issues or concerns. Second are ML detection models; these take longer to update and require additional validation due to their potential to cause unexpected shifts in the detection distribution. Third, the orchestration logic that integrates the previous components into the final detection (Fig. LABEL:fig:production).
[…]
fig:production Refer to caption
@@ 4.1 Signature Updates
4.2 Model and Logic Updates
@@
5 Previous Work
As discussed earlier in this paper, significant research has focused on developing detection methods for LLM attacks and protections against them, consisting of technical controls and operational best practices OWASP Foundation (2025); Jiang et al. (2023); Christiano (2022); Bullwinkel et al. (2025); Jain et al. (2023). But relatively few works address the operational challenges of deploying these defenses in production environments requiring rapid threat response and continuous adaptation. We have not found similar examples of a dedicated threat intelligence function designed specifically for AI-related threats that considers the expanded attack surface and the deepening integrations of generative AI and agentic AI tools in software systems and technology.
Early work in operationalizing LLM security began with the development of Vigil Swanda (2023) in 2022, which introduced the application of YARA rules to the domain of LLM threat detection. The open-source project demonstrated that pattern-matching techniques could effectively identify prompt injection attempts and other LLM-specific threats while leveraging existing security tooling rather than requiring specialized development of novel tools. Our comprehensive defense platform described in this paper significantly extends this foundational approach by integrating rule-based detection into an enterprise-grade threat intelligence pipeline with automated deployment capabilities, enabling organization-scale protection with the ability to rapidly respond to emerging threats.
Recent advances in prompt injection detection have produced promising approaches, including research that compares the performance of early detection systems for LLM security Gakh and Bahsi (2025). Google DeepMind’s CaMeL Debenedetti et al. (2025) takes an architectural approach, applying traditional software security principles like control flow integrity to LLM systems. PromptShield Jacob et al. (2025) achieves detection with low false positive rates through fine-tuning detectors on both conversational and application-structured data, while Attention Tracker Hung et al. (2024) introduces training-free detection by analyzing attention patterns within LLMs. However, Hackett et al. Hackett et al. (2025) demonstrated that many prominent guardrail systems can be bypassed with up to 100% success rates using character injection techniques, highlighting the limitations of point-in-time detection approaches. While comprehensive prevention remains an open problem, our platform reduces the window of vulnerability by translating new threats into protections quickly through rapid signature updates and safe staged deployments.
6 Conclusion